# Unsettling the Hegemony of Intention: Agonistic Image Generation

### v2

@@ 3. Four Paradigmatic Interfaces for Image Generation Interactions
Figure 1. Annotated screenshots from each of the paradigmatic image generation interfaces evaluated in our study. See §3 for details.
@@
To holistically evaluate how agonistic design might encourages reflection (§2.4), we create four image generation interfaces: Baseline, which accepts the user’s prompt and outputs generated images with no further interaction, and three paradigmatic interfaces each embodying a dominant approach or value — Diverse for diversity, Reformulative for intention actualization, and Agonistic for agonism. In this section, we describe the important frontend and backend elements of each interface, with more details in Appendix §B. Annotated screenshots are displayed in Figure 1.
@@
Baseline  $\cdot$  We design a baseline interface as a control for the study. In this interface, users enter a prompt and are displayed four images generated from the prompt. We use a lightweight open-source image generation model (Black Forest Labs, flux-schnell) throughout all interfaces in the study, to avoid relying on interfaces that perform under-the-hood prompt rewriting (like the DALL-E API).
@@
Diverse  $\cdot$  Diverse represents the paradigm of image generation that explicitly corrects for diversity issues by rewriting prompts in more diverse ways, as was reported in the Gemini case. To create this interface, we use a leaked alleged DALL-E prompt as reported by (Milmo and Kern, 2024), which instructs the model to “diversify depictions of ALL images with people to include DESCENT and GENDER for EACH person using direct term.” We instruct GPT-3.5 to rewrite the user prompt with the above instructions and use the rewritten prompt to generate four images that are then displayed to the user. Although we recognize that the prompt may not match the exact approach used in the Gemini case, it anecdotally yields similar results to reported images for a variety of historical prompts. Moreover, this approach makes Diverse a realistic interface to compare against, since it uses the (alleged) prompt for the DALL-E API, which is widely in use.
@@
Reformulative  $\cdot$  The Reformulative interface embodies the value of intention actualization and is inspired by §2.3; it supports users by producing prompt reformulations that add detail to the prompt (e.g., adding useful semantics or stylistic indicators) in ways that are likely to result in more aesthetic or preferred images. This interface is similar to interfaces like Promptify, which allows users to iteratively refine their prompts based on a variety of AI-generated reformulations (Brade et al., 2023). Skilton and Cardinal propose using AI to generate a list of “suggested descriptors” that the user can then choose from and modify to create images that depart from stereotypical representations (Skilton and Cardinal, 2024). We instruct GPT-4o to generate a diverse set of detail-adding reformulations of the user prompt with in-context examples from or inspired by Brade et al. To aid visual navigation of reformulations, we generate a sample image (“thumbnail”) for each reformulation and display both in a list of suggested reformulations to the user. The user can choose a suggestion and modify the suggestion before using it to generate images. This interface is designed to have a comparable interaction level and similar features where applicable to Agonistic for fair comparison.
Agonistic  $\cdot$  The Agonistic interface attempts to forefront ambiguities and controversies over visual interpretation to the user, drawing from §2.2 and §2.4. We implement Agonistic by creating a multi-step retrieval-augmented workflow that researches controversies about the user prompt from Wikipedia and presents them to the user. Our work builds on insights from Cox et. al., who find that presenting users with maximally distant phrases (measured by language embeddings) is effective at enhancing diversity in creative ideation (Cox et al., 2021). Because Wikipedia allows public contributions, we believe it is more likely to capture relevant prompt controversies. $\cdot$  We identify the main subject of the prompt with GPT-3.5 and perform a Wikipedia search for 50 pages related to the main subject. We then use GPT-4o to filter for 40 pages relevant to the user prompt based on their titles (excluding, for example, soccer player Gabriel Jesus from a search for “Jesus”). Next, we compute a controversy score for each page and rank pages by controversy, selecting the top 20 most controversial pages. The controversy score is calculated by dividing the total number of edits by the number of unique editors, following findings from Kittur et. al. that controversy is positively correlated with total number of edits and negatively correlated with the number of unique editors (Kittur et al., 2007). We chose this particular controversy metric for its efficiency and strong qualitative performance.  $\cdot$  We then instruct GPT-4o to generate 5 mental images an “average person” might have of the main subject and provide this in-context to the interpretation generation call. For each page, GPT-4o then selects 4 sections based on their titles to read and produces an interpretation with 4 fields: 1) a section summary explaining the main points of the cited page section; 2) a description of what the user’s main subject looks like, taking into account the user’s full prompt; 3) a source with the referenced page and section; 4) a justification explaining how the section content justifies the description. The section summary is not shown to the user but is generated first to reduce model hallucinations.  $\cdot$  We instruct GPT-4o to generate descriptions that challenge the provided mental images, and phrase justifications in the form “you may assume…, but… .” These decisions were made to increase the likelihood that users would encounter discomforting suggestions, creating opportunities for critical reflection. Finally, like in Reformulative, we generate a “thumbnail” for each interpretation. Users are shown a list of interpretations with descriptions, thumbnails, and sources; when users click on an interpretation, the card expands and the justification becomes visible. After a 3-second wait period to encourage users to read the justification, the user can “accept” the interpretation and generate images, as well as edit the description text and re-generate. The design iteration of this interface is documented in §B.1.
@@ 4. Experimental Design
To evaluate all four interfaces with respect to reflection, we conduct a within-subjects comparative lab study and ask (RQ1): How each interface induces reflection using a variety of measures for user reflection; and (RQ2): Why each interface induces reflection by examining how other variables could explain the reflection observations in RQ1.
@@ 4.1. User Task and Interview Structure
Setup ($\sim$15 mins). Each participant was randomly assigned to an ordering of interfaces, with Baseline fixed as the first interface (e.g., [Baseline Agonistic, Diverse, Reformulative] — 6 unique orderings), to account for learning effects between non-Baseline interfaces. The interviewer received consent to record the interview and use anonymized data. Then, to familiarize the participants with the interfaces, participants were guided through using each interface in their assigned order with the prompt “a person.” Afterwards, participants selected a subject (prompt) for their task; they were told to choose one of interest to them and encouraged to choose one in a randomly pre-assigned category. Subjects all featured people and were divided into three categories — identity / demographics, history, and political issues. This choice of categories allowed for potential reflection on issues of visual representation. Participants were provided with examples in each category for inspiration but also allowed come up with their own subjects. See §A.1 for participants’ chosen prompts and §A.2 for example prompts.
User Task ($\sim$45 mins). After setup, participants were informed of their task: to create a collage of ten images which represented all relevant aspects of the subject. This task forced participants to make nuanced choices about what to include and exclude in visual representation of their chosen subject. Participants constructed an initial collage of ten images with Baseline. Then, they improved their collage by interacting with the other interfaces (Diverse, Reformulative, Agonistic) in their assigned order. If participants produced an image they wanted to add to their collage, they picked one of their existing images to replace. The task design of accumulating collage improvement gave permanence to users’ decisions about visual representation, provided a holistic picture of how users interacted with interfaces across different stages of exploration, and was a more natural and engaging task for participants compared with other experiment designs tested. Participants were instructed to “think out loud,” commenting on their choices to include or exclude images in their collage. Figure 8 displays an example collage progression from the study.
@@ 4.3. Measures
Given the work discussed in §2.4, we measured reflection by looking quantitative and qualitative signals of how much individuals questioned or changed prior assumptions after using our interfaces. While it is more common to use purely qualitative metrics when measuring reflection (Bentvelzen et al., 2022), we believed a mixed-methods approach offered unique benefits for our study because quantitative metrics would allow us to more precisely compare reflection across interfaces. We drew from common metrics in mixed methods studies of reflection such as interviews, questionnaires, and self-reports (Bentvelzen et al., 2022).
@@
Interview Coding. After conducing interviews, the two co-first authors coded interview transcripts to systematically extract further information about user experience. They independently coded a small subset of interviews before coming together to build a coding ontology. We divided codes into intents (how user intents changed while adding an image, roughly correlated to different degrees of reflection) and values (reasons users give for adding an image). Values are not mutually exclusive, because a user can invoke multiple values when adding an image. The two first co-authors then independently coded a subset of three interviews to compute inter-rater reliability (IRR), revising the coding ontology and independently re-coding interviews as needed to resolve disagreements. The final IRR was 0.67 based on a weighted average of Cohen’s Kappa scores across value codes. After coding all interviews, the two co-first authors reviewed all images marked with non-Direct intent together to establish more agreement on intent codes. See §C for our IRR methodology and §5.2, §6.2 for a delination of intents and values.
@@ 5. How do interfaces induce reflection?
(a) Overall, by interface.
(b) Breakdown within interface by prompt categories: “H” for History, “P” for Politics, “I” for identity.
Figure 2. Participant ratings on a 5-point scale for how much interacting with the interface made them rethink their mental picture.
@@
In this section, we explore how different interfaces induce reflection by investigating variables that directly measure some aspect of reflection collected in the study. The aim is to paint a holistic picture of how reflection occurred under different interfaces. In Table 1 we provide two examples of reflection for each of the interfaces from our interviews. To characterize reflection across different interfaces, we investigate users’ self-reported change in mental image (§5.1) and reflection on a per-image level (§5.2) as described in §4.3.
PID Examples of Participant Reflection
Baseline P24 The participant noticed that all the images for “the constitutional convention” featured the delegates sitting. In subsequent generations, they specified that the delegates should be standing. They reflected: “when I kept generating those prompts and they were sitting […] I don’t think I would have specified whether they were sitting or standing before. Since I thought about it, I had to… play around with the prompt to get them to stand”. They further remarked: “Interacting with this interface […] made me pay more detailed attention to what I perceive would have actually happened at that historical event.
P4 After generating a few images for the subject “a queer community” which seemed to “rely on heavy stereotypes […] like pride parades”, the participant wanted to generate “everyday examples of queer people”. After producing a few images, they remarked “now you lose the queer part.” After a few moments, they reflected: “Well now you really got me questioning myself […] what is a queer community? Do they all have to be wearing pride flags? How do you tell? Now I’m getting in my own head about it.
Diverse  P19 The participant was attempting to generate an image of “A Syrian refugee working in Germany” when Diverse generated an image of a woman in a healthcare setting. The participant paused and remarked, “can you train to be a doctor [in] 12 years?” They later explained that they were unsure whether a Syrian refugee would have been able to study long enough to become a licensed medical professional since leaving Syria, but reasoned that “the situation has gone on long enough [that] they have studied and integrated in the societies they’re in, where [they’re] able to actually finish something in the medical sciences.
P29 The participant had previously generated all images of White people for the prompt “a Jewish person” using Baseline. After Diverse produced images of Black people, the participant felt that the images were not appropriate but became confused upon further reflection: “Okay, I’m focused on the Black […] person here because, this is because my lack of knowledge […] I don’t know what makes a person Jewish other than the religious element [but…] to be Jewish, I think you have to be born in a family […] I could be wrong.
Reformulative P10 With the prompt “A gun owner,” Reformulative generated the suggestion, “An elderly man polishing his grandfather’s antique shotgun.” Upon reading the suggestion, the participant reacted, “I didn’t even think about that kind [of] collector gun owner,” before using the suggestion to generate and add an image to their collage. They later reflected that the interface “reminded [them] of things [they] forgot about” and “sparked a new mental image” for them.
P19 Given the prompt “A gun control advocate,” Reformulative generated a suggestion of “A gun control advocate engaging in a community discussion.” The participant reacted, “I like that it’s suggested positioning the advocate in a community space, where all of the other images [from previous interfaces] were very portrait-style.” They used the suggestion to generate and add an image to their collage, later sharing that they liked how Reformulative helped them “think of things that [they] wouldn’t think to ask.
Agonistic P5 After interacting with Agonistic on the subject “World War II” which showed many less well-known and more local actualizations of WWII, the participant (whose family was strongly impacted by WWII) revised their design statement from “WWII involved mass conflicts that…” to “WWII involved conflicts that…”. They reflected: “I think it just involves conflicts in general now. It’s making me rethink that word [mass]: history tends to focus on real large-scale things, but I find that all of them are important to me personally. I feel like it’s actually made me reconnect to the way I think about World War II. And instead of trying to project what I would like, think that other people need to see about World War II.
P28 The participant initially described their mental image of the prompt “the signing of the Declaration of Independence” as “A group of […] white men like Thomas Jefferson standing around a table signing a document.” When Agonistic generated images of non-White people, drawing from sources like “Haitian Declaration of Independence,” the participant observed, ”when […] I put Declaration of Independence, I assume a mental image of the US Declaration of Independence, but so many nations have Declarations of Independence.” They continued, “It makes me realize […] this assumption of bias… Oh my gosh, I was so narrow-minded in my approach.
Table 1. Two examples of reflection observed during our interviews when using each interface.
5.1. Change in mental image
The change in mental image is reported by participants after they use each interface and is measured by agreement with the statement “Interacting with the interface made me rethink the mental picture of the subject I had right before using this interface” on a 5-point Likert-style scale. We take higher reported changes in mental image to be indicative of more reflection. Averages are reported in Figure 2 and Table 4.
Agonistic induces the most reflection as measured by self-reported rethinking. Participants’ mental images change more under Agonistic than Reformulative ($p\ll 0.01$), and more under Reformulative than Diverse ($p=0.05$). Neither Reformulative nor Diverse surpass the Baseline(likely due to the fact that participants interact with Baseline first), but Agonistic does ($p\ll 0.01$). The ‘hierarchy’ of interfaces by reflection is thus $\left[\textsf{{Agonistic}}>\{\textsf{{Baseline}},\textsf{{Reformulative}}\}>\textsf{{Diverse}}\right]$. A similar ‘hierarchy’ emerges when considering how participants rank interfaces by how much their mental image changed retroactively at the end of the study (see §D.1), as well as measuring how much participants’ “design statements” changed after using each interface (see §D.3). Agonistic further induces greater unique rethinking based on an analysis of ordering effects (see §D.2).
Rethinking under Agonistic is robust to prompt category. Though Diverse and Reformulative see significant drops in rethinking for historical prompts compared to political prompts ($-0.27$ and $-1.27$, respectively) the same drop is not observed for Agonistic ($-0.00$). While all domains have roughly similar rethinking under Agonistic, political subjects lend themselves to more rethinking overall under Diverse and Reformulative than historical subjects.
5.2. Reflection when adding images
Next, we aim to capture a richer picture of reflection by analyzing the distribution of intent codes across interfaces. As explained in §4.3, intents code for how user intents changed while adding an image to their collage. Intent codes are defined as follows: 1) Direct: the user already intended to generate the image; 2) Reminder: the user is reminded of a detail about the prompt they already accepted but forgot; 3) Expansion: the user accepts an image as a different valid interpretation of the prompt than their original intent; 4) Challenge: the user realizes an error or otherwise experiences a significant change in their original intent. Direct intent represents little to no reflection, whereas Reminder, Expansion, and Challenge represent increasing degrees of reflection. The distribution of intents across interfaces is shown in Fig. 4
Agonistic induces the most reflection as measured by quantitative interview analysis. Participants replace the most images on average with Agonistic(3.1), followed by Reformulative (2.9) and Diverse (1.7). We observe Direct intent most frequently with Baseline ($98\%$), followed by Diverse ($95\%$), Reformulative ($90\%$), and a sharp drop in Agonistic ($67\%$). Agonistic also has the highest proportion of Reminder and Expansion of intent across all interfaces ($11\%$ and $19\%$ respectively, compared to the next-highest $4\%$ and $6\%$ for Reformulative), and is the only interface to Challenge  intent ($4\%$). An example of Challenge intent by P28 is given in Table 1. A minor difference with §5.1 is the finding that Diverse induces more frequent reflection than Baseline (see §D.6).
Figure 3. Breakdown of how user intents changed when adding images for each interface (discussed in‘§5.2; Table 7).
@@ 5.2. Reflection when adding images
Figure 4. Mean responses on a 5-point scale for Rethink, Appropriateness, Control, and Interest (discussed in §6.1; Table 5).
6. Why do different interfaces induce reflection?

### v3

@@ 3. Four Image Generation Interfaces
Figure 1. Screenshots from each image generation interface evaluated in lab study. See §3 for details. Note that  Agonistic  depicts men of color as a possible interpretation of “Founding Father,” drawing upon the play Hamilton. This interpretation is displayed among many others to engage users in discourse about what a Founding Father looks like.
@@
To holistically evaluate how agonistic design affects user reflection, we create a Baseline interface, which accepts the user’s prompt and outputs generated images with no further interaction, and three paradigmatic interfaces each embodying a dominant approach or value: Diverse for diversity, Reformulative for intention actualization, and Agonistic for agonism. In this section, we describe the key frontend and backend elements of each interface.
@@
Baseline  $\cdot$  We design a Baseline interface as a control for the study. In this interface, users enter a prompt and are displayed four images generated from the prompt. We use a lightweight open-source image generation model (Black Forest Labs, flux-schnell) throughout all interfaces in the study, to avoid relying on interfaces that perform under-the-hood prompt rewriting (e.g., DALL-E API).
@@
Diverse  $\cdot$  The Diverse interface represents a paradigm of image generation that explicitly corrects for diversity issues by rewriting prompts in more diverse ways, as reported in the Gemini case. To create this interface, we use an allegedly leaked DALL-E prompt as reported by Milmo and Kern (2024), which instructs the model to “diversify depictions of ALL images with people to include DESCENT and GENDER for EACH person using direct term.” We instruct GPT-3.5 to rewrite the user prompt with the above instructions and use the rewritten prompt to generate four images that are then displayed to the user. Although we recognize that the prompt may not match the exact approach used in the Gemini case, it anecdotally yields similar results to reported images for a variety of prompts.
@@
Reformulative  $\cdot$  The Reformulative interface embodies the value of intention actualization and is inspired by §2.3; it supports users by producing prompt reformulations that add detail to the prompt (e.g., adding useful semantics or stylistic indicators) in ways that are likely to result in more aesthetic or preferred images. This interface is similar to interfaces like Promptify, which allows users to iteratively refine their prompts based on a variety of AI-generated reformulations (Brade et al., 2023). Similarly, Skilton and Cardinal (2024) propose using AI to generate a list of “suggested descriptors” that the user can then choose from and modify to create images that depart from stereotypical representations. We instruct GPT-4o to generate a diverse set of detail-adding reformulations of the user prompt with in-context examples from or inspired by Promptify. To aid visual navigation of reformulations, we generate a sample image (“thumbnail”) for each reformulation and display both in a list of suggested reformulations to the user. The user can choose a suggestion and modify the suggestion before using it to generate images. This interface is designed to have a comparable interaction level and similar features to Agonistic for fair comparison.
Agonistic  $\cdot$  The Agonistic interface aims to provide resources for users to engage in political controversies over visual representations of their prompt. Drawing from §2.2 and §2.4, we take inspiration from Angeli et al. (2021) to derive the following tenets of agonistic design for image generation tools: (T1): confront users with unsettling content that challenges user assumptions; (T2): encourage engagement with relevant ongoing political controversies; (T3): respect users as adversaries by allowing them to navigate conflicts on their own terms.
We implement Agonistic with a multi-step retrieval-augmented workflow that researches controversies about the user prompt from Wikipedia and presents them to the user (Fig. 2). We identify the main subject of the prompt with GPT-3.5, perform a Wikipedia search for 50 pages related to the main subject, and use GPT-4o to filter for the 40 pages most relevant to the user prompt by title. Next, we compute a controversy score for each page by dividing the number of edits by the number of unique editors, following findings from Kittur et al. (2007) that controversy is positively correlated with number of edits and negatively correlated with number of unique editors. We choose this particular controversy metric for its efficiency and strong qualitative performance in preliminary tests. Finally, we rank the pages by controversy and select a random subset of 10 from the top 20 most controversial pages. We select Wikipedia as a research source following because its public editing process lends itself well to capturing the types of public controversies characteristic of agonistic pluralism, as suggested by as Crawford (2016) (T2).
Using this set of 10 pages, we generate interpretations of the user’s prompt. We first instruct GPT-4o to generate 5 mental images an “average person” might have of the main subject. For each page, we instruct GPT-4o to generate an interpretation with 4 fields: 1) a section summary explaining the main points of the referenced page section; 2) a visual description of the prompt based on information from the section; 3) a source with the referenced page and section; 4) a justification explaining how the section content justifies the description. We instruct GPT-4o to generate descriptions the challenge the provided mental images to increase the likelihood that users encounter discomforting suggestions (T1). The section summary is not shown to the user but is generated first to reduce model hallucinations. Finally, like in Reformulative, we generate a “thumbnail” for each interpretation. Users are shown a list of interpretations with descriptions, thumbnails, and sources. When users click on an interpretation, the card expands to show the justification. After a 3-second wait period to encourage users to read the justification, the user can “accept” the interpretation and generate images, as well as edit the description text and re-generate (T3).
Figure 2. Sample agonistic interface generation workflow.
@@ 4. Experimental Design
To evaluate how each interface affects user reflection, we conducted a within-subjects comparative lab study and asked (RQ1): How each interface induces reflection using a variety of measures for user reflection; and (RQ2): Why each interface induces reflection by examining how other variables could explain the reflection observations in RQ1.
@@ 4.1. User Task and Interview Structure
Setup ($\sim$15 mins). The interviewer received consent to record the interview and use anonymized data. Each participant was assigned to a random ordering of interfaces, with Baseline fixed as the first interface, to account for learning effects between non-Baseline interfaces. Then, to familiarize the participants with the interfaces, participants were guided through using each interface in their assigned order with the prompt “a person.” Afterwards, participants selected a subject (prompt) for their task; they were told to choose one of interest to them and encouraged to choose one in a randomly pre-assigned category. Subjects all featured people and were divided into three categories: identity/demographics, history, and politics. This choice of categories allowed for potential reflection on issues of visual representation. Participants were provided with examples in each category for inspiration but also allowed come up with their own subjects. See §A.1 for participants’ chosen prompts and §A.2 for example prompts.
User Task ($\sim$45 mins). After setup, participants were informed of their task: to create a collage of ten images which represented all relevant aspects of the subject. This task forced participants to make nuanced choices about what to include and exclude in visual representation of their chosen subject. Participants constructed an initial collage of ten images with Baseline. Then, they improved their collage by interacting with the other interfaces (Diverse, Reformulative, Agonistic) in a preassigned order. If participants produced an image they wanted to add to their collage, they picked one of their existing images to replace. The task design of accumulating collage improvement gave permanence to users’ decisions about visual representation, provided a holistic picture of how users interacted with interfaces across different stages of exploration, and was a more natural and engaging task for participants compared with other experiment designs tested. Participants were instructed to “think out loud,” commenting on their choices to include or exclude images in their collage. Fig. 9 displays an example collage progression from the study.
@@ 4.3. Measures
Given the definition of reflection discussed in §2.4, we measured reflection by looking quantitative and qualitative measures of how much individuals questioned or changed prior assumptions after using our interfaces. While it is more common to use purely qualitative metrics when measuring reflection (Bentvelzen et al., 2022), we believed a mixed-methods approach offered unique benefits for our study because quantitative metrics would allow us to more richly compare reflection across interfaces. Because of our small sample size, the quantitative methods were selected for the purpose of making broad observations about induced reflection rather than precise claims. We drew from common metrics in mixed methods studies of reflection such as interviews, questionnaires, and self-reports (Bentvelzen et al., 2022).
@@
Interview Coding. After conducting interviews, the two co-first authors coded interview recordings to systematically extract further information about user experience. They independently coded a small subset of interviews before coming together to build a coding ontology. We divided codes into intents (how user intents changed while adding an image, roughly correlated to different degrees of reflection) and values (reasons users give for adding an image). Values are not mutually exclusive, because a user can invoke multiple values when adding an image. Interviews were coded on the level of individual images added: for each added image, the authors recorded which intent and value codes were invoked by the user. The two first co-authors then independently coded a subset of three interviews to compute inter-rater reliability (IRR), revising the coding ontology and independently re-coding interviews as needed to resolve disagreements. After coding all interviews, the two co-first authors further reviewed all images marked with non-Direct intent together to establish higher agreement on intent codes. The final IRR was 0.67 based on a weighted average of Cohen’s Kappa scores across value codes. See §5.2 and §6.2 for a taxonomy of intent and value codes, and §C for the IRR methodology.
Baseline After observing that all the images for “the constitutional convention” featured the delegates sitting, the participant specified that the delegates should be standing. They reflected: “when I kept generating those prompts and they were sitting […] I don’t think I would have specified whether they were sitting or standing before. Since I thought about it, I had to… play around with the prompt to get them to stand”. They further remarked: “Interacting with this interface […] made me pay more detailed attention to what I perceive would have actually happened at that historical event.” (P24)
After generating images for the subject “a queer community” which seemed to “rely on heavy stereotypes […] like pride parades”, the participant wanted to generate “everyday examples of queer people”. After producing a few images, they remarked “now you lose the queer part.” They reflected: “Well now you really got me questioning myself […] what is a queer community? Do they all have to be wearing pride flags? How do you tell?” (P4)
Diverse  The participant was attempting to generate an image of “A Syrian refugee working in Germany” when Diverse generated an image of a woman in a healthcare setting. The participant paused and remarked, “can you train to be a doctor [in] 12 years?” They later explained that they were unsure whether a Syrian refugee would have been able to study long enough to become a licensed medical professional since leaving Syria, but reasoned that “the situation has gone on long enough [that] they have studied and integrated in the societies they’re in, where [they’re] able to actually finish something in the medical sciences.” (P19)
The participant had previously generated all images of White people for the prompt “a Jewish person” using Baseline. After Diverse produced images of Black people, the participant felt that the images were not appropriate but became confused upon further reflection: “Okay, I’m focused on the Black […] person here because, this is because my lack of knowledge […] I don’t know what makes a person Jewish other than the religious element [but…] to be Jewish, I think you have to be born in a family […] I could be wrong.” (P29)
Reformulative With the prompt “A gun owner,” Reformulative generated the suggestion, “An elderly man polishing his grandfather’s antique shotgun.” Upon reading the suggestion, the participant reacted, “I didn’t even think about that kind [of] collector gun owner,” before using the suggestion to add an image to their collage, later reflecting that the interface “reminded [them] of things [they] forgot about” and “sparked a new mental image”. (P10)
Given the prompt “A gun control advocate,” Reformulative generated a suggestion of “A gun control advocate engaging in a community discussion.” The participant reacted, “I like that it’s suggested positioning the advocate in a community space, where all of the other images [from previous interfaces] were very portrait-style.” They used the suggestion to generate and add an image to their collage, later sharing that they liked how Reformulative helped them “think of things that [they] wouldn’t think to ask.” (P19)
Agonistic After interacting with Agonistic on the subject “World War II” which showed many less well-known and more local actualizations of WWII, the participant (whose family was strongly impacted by WWII) revised their design statement from “WWII involved mass conflicts that…” to “WWII involved conflicts that…”. They reflected: “I think it just involves conflicts in general now. It’s making me rethink that word [mass]: history tends to focus on real large-scale things, but I find that all of them are important to me personally. I feel like it’s actually made me reconnect to the way I think about World War II. And instead of trying to project what I would like, think that other people need to see about World War II.” (P5)
The participant initially described their mental image of the prompt “the signing of the Declaration of Independence” as “A group of […] white men like Thomas Jefferson standing around a table signing a document.” When Agonistic generated images of non-White people, drawing from sources like “Haitian Declaration of Independence,” the participant observed, ”when […] I put Declaration of Independence, I assume a mental image of the US Declaration of Independence, but so many nations have Declarations of Independence.” They continued, “It makes me realize […] this assumption of bias… Oh my gosh, I was so narrow-minded.” (P28)
Table 1. Two examples of reflection observed during lab study for each interface.
@@ 5. How do interfaces induce reflection?
In this section, we aim to paint a holistic picture of how reflection occurred under different interfaces by reporting results that directly measure some aspect of reflection collected in the study, including users’ self-reported change in mental image (§5.1) and reflection on a per-image level (§5.2).
@@
Figure 3. Mean responses on a 5-point scale for Rethink, Appropriateness, Control, and Interest (§5.1, 6.1; Table 4).
Figure 4. Breakdown of reported rethinking by prompt category: History (‘H’), Politics (‘P’), Identity (‘I’) (§5.1; Table 5).
5.1. Change in Mental Picture
Participants self-reported change in mental picture after using each interface by rating agreement with the statement “Interacting with the interface made me rethink the mental picture of the subject I had right before using this interface” on a 5-point Likert-style scale. We interpret higher reported scores as indicative of more reflection.
Agonistic induces the most reflection as measured by self-reported rethinking (Fig. 3). Participants report greater rethinking under Agonistic than Reformulative ($p\ll 0.01$), and under Reformulative than Diverse ($p=0.05$). Neither Reformulative nor Diverse surpass Baseline (likely because participants interacted with Baseline first), but Agonistic does ($p\ll 0.01$). The ‘hierarchy’ of interfaces in terms of self-reported reflection is thus $\left[\textsf{{Agonistic}}>\{\textsf{{Baseline}},\textsf{{Reformulative}}\}>\textsf{{Diverse}}\right]$.
Rethinking under Agonistic is robust to prompt category (Fig. 4). Though Diverse and Reformulative see significant drops in rethinking for historical prompts compared to political prompts ($-0.27$ and $-1.27$, respectively) the same drop is not observed for Agonistic ($-0.00$). For historical topics, rethinking is low for Diverse (1.33) and Reformulative (1.33), but much higher for Agonistic (3.00). For political topics, rethinking is high for both Agonistic (3.00) and Reformulative (2.60). While all prompt categories have roughly similar rethinking under Agonistic, political subjects lend themselves to more rethinking overall under Diverse and Reformulative than historical subjects.
5.2. Reflection Induced by Adding Images
Figure 5. Breakdown of intent codes per interface (§5.2; Table 6).
Next, we analyze the distribution of “intent” codes across interfaces. As explained in §4.3, “intents” code for how user intents changed while adding an image to their collage. We define the following intent codes: 1) Direct: the user already intended to generate the image; 2) Reminder: the user is reminded of an aspect the prompt they already accepted but forgot; 3) Expansion: the user accepts an image as a different valid interpretation of the prompt than their original intent; 4) Challenge: the user realizes an error or otherwise experiences a significant change in their original intent. Direct intent represents little to no reflection, whereas Reminder, Expansion, and Challenge represent increasing degrees of reflection.
[…]
6. Why do interfaces induce reflection?